# Adversarial Inception for Bounded Backdoor Poisoning in Deep Reinforcement Learning

### v1

@@ Abstract
Recent works have demonstrated the vulnerability of Deep Reinforcement Learning (DRL) algorithms against training-time, backdoor poisoning attacks. These attacks induce pre-determined, adversarial behavior in the agent upon observing a fixed trigger during deployment while allowing the agent to solve its intended task during training. Prior attacks rely on arbitrarily large perturbations to the agent’s rewards to achieve both of these objectives - leaving them open to detection. Thus, in this work, we propose a new class of backdoor attacks against DRL which achieve state of the art performance while minimally altering the agent’s rewards. These “inception” attacks train the agent to associate the targeted adversarial behavior with high returns by inducing a disjunction between the agent’s chosen action and the true action executed in the environment during training. We formally define these attacks and prove they can achieve both adversarial objectives. We then devise an online inception attack which significantly out-performs prior attacks under bounded reward constraints.
@@ 1 Introduction
Reinforcement learning (RL) algorithms are versatile tools allowing artificial agents to optimize complex tasks directly from interactions with their environment. The most popular and powerful RL approaches, like Proximal Policy Optimization (PPO) (Schulman et al., 2017) and Deep Q-Networks (DQN) (Mnih et al., 2013), utilize deep-neural networks as function approximators, forming an area of research referred to as Deep Reinforcement Learning (DRL). This versatility has lead to the ever growing adoption of DRL based approaches in safety and security critical domains, such as automated cyber defenses (Vyas et al., 2023), self-driving vehicles (Kiran et al., 2021), robotic warehouse management (Krnjaic et al., 2023), and space traffic coordination (Dolan et al., 2023).
The wide-spread applicability of DRL makes it a target for external adversaries wishing to influence an agent’s behavior. This necessitates deeper studies into the capabilities of adversarial attacks against DRL so that practitioners can know how to defend against them. Thus, in this work we focus our efforts towards a better understanding of backdoor poisoning attacks which manipulate the training of an agent such that their behavior can be directly controlled during deployment upon observing a pre-determined “trigger”. Multiple works (Kiourti et al., 2019; Cui et al., 2023; Rathbun et al., 2024; Wang et al., 2021) study these attacks assuming the adversary can arbitrarily alter the agent’s rewards. We demonstrate how defenders can easily detect these attacks during training without any strong assumptions about the underlying task. We then show how clipping the adversary’s reward perturbations to evade detection severely diminishes these attacks’ capabilities.
These shortcomings necessitate the development of a new backdoor poisoning approach capable of guaranteeing the adversary’s success while operating under a restricted reward poisoning setting. To this end we provide multiple contributions towards answering the question: Can backdoor attacks be successful in DRL without arbitrarily manipulating rewards? Specifically we:
@@
    Highlight the detectability and theoretical limitations of prior attacks in the bounded reward poisoning setting using intuitive examples.
@@
    Formulate the adversarial inception attack framework (visualized in Figure 1) which uses novel action manipulation techniques to guarantee backdoor attack success while maintaining the victim agent’s performance in their intended task.
@@
    Develop a novel backdoor poisoning attack “Q-Incept” leveraging adversarial inception and DQN based techniques to achieve significant increases in attack success over prior attacks.
@@
    Provide in-depth evaluation of Q-Incept on environments spanning self driving, video game playing, cyber network defending, and robotic navigating tasks.
Figure 1: Visualization of inception attacks. During training the agent observes the trigger and chooses the target action (Turn Right), but transitions with respect to the optimal action (Forward). After deployment this relationship no longer exists, causing the agent to drive off the road instead.
@@ 2 Related Work and Background
Attack Q-Incept (Ours) SleeperNets TrojDRL BadRL
Reward Poisoning Bounded Unbounded Static Static
Action Manipulation Inception None Forced Forced
Attack Success Guarantees of Existing Backdoor Attacks in DRL
Unbounded Rewards Yes Yes No No
Bounded Rewards Yes No No No
Table 1: Attack success guarantees of different backdoor attacks including our proposed Q-Incept.
Here we provide an overview of the existing backdoor attack literature in DRL. When executing a backdoor attack the adversary manipulates the Markov Decision Process (MDP) which an agent is being trained to optimize. These MDPs are often defined as $M=(S,A,R,T,\gamma)$ where $S$ is the set of states in the environment, $A$ is the set of possible actions for the agent to take, $R:S\times A\times S\rightarrow\mathbb{R}$ is the reward function, $T:S\times A\times S\rightarrow[0,1]$ represents the transition probabilities between states given actions, and $\gamma\in[0,1]$ is the discount factor. All the attacks discussed in this paper, including our own, are designed to target agents trained using DRL algorithms like PPO or DQN.
Backdoor attacks against DRL were first explored by Kiourti et al. (2019) whose attack, TrojDRL, showed success against DRL agents training on Atari games (Brockman et al., 2016). Since then multiple other works (Wang et al., 2021; Yang et al., 2019; Yu et al., 2022; Cui et al., 2023) have used similar approaches in different domains – all statically altering the agent’s reward to some fixed ±$\pm c$, and many forcing the agent to take the targeted action $a^{+}$ during training.
While effective in some domains, Rathbun et al. (2024) proved these static reward poisoning approaches are insufficient for achieving the adversary’s objectives. This motivated their development of a new, unbounded reward poisoning attack, SleeperNets, with strong guarantees of attack success. In Table 1 we summarize these guarantees and highlight, in bold, new theoretical results developed in our work. In spite of its strong theory, SleeperNets suffers from the arbitrarily large perturbations it applies to the agent’s reward. This leaves the attack open to detection by a defender monitoring the agent’s reward signal for outliers, which is explored in Section 3.2. Other works (Franzmeyer et al., 2022; Nasvytis et al., 2024; Bharti et al., 2022) have explored the test-time detectability of adversarial attacks assuming access to a clean set of environment interactions. In this work we explore detection techniques making no such assumptions, allowing a defender to detect prior attacks during training using only limited knowledge about the benign MDP’s reward function.
@@ 3 Problem Formulation
In backdoor attacks against DRL there are two primary parties. The victim attempts to train an agent on benign MDP $M=(S,A,R,T,\gamma)$ with some stochastic algorithm $\mathcal{L}(M)$ returning policy $\pi:\mathbb{S}\times A\rightarrow[0,1]$. The adversary attempts to instead train the agent on adversarial MDP $M^{\prime}=(S^{\prime},A,R^{\prime},T^{\prime},\gamma)$ with the goal of inducing adversarial behavior in the agent upon observing a pre-determined trigger. We model the trigger as a function $\delta:S\rightarrow\mathbb{S}$ which applies a fixed pattern to a given input state. Here $\mathbb{S}$ represents the space within which $S\subseteq\mathbb{S}$ resides, and $S_{p}\doteq\text{im}(\delta(S))$ refers to the set of poisoned states. In the next subsection we will define our objectives adapted from Rathbun et al. (2024) and Kiourti et al. (2019). Then in the following subsection we will show how this formulation leads to trivially detectable attacks, motivating additional problem constraints.
@@ 3.1 Adversarial Objectives
In this work, we will be implementing “targeted attacks” (Kiourti et al., 2019) in which the desired adversarial behavior is a fixed action $a^{+}\in A$. This objective is the current standard for backdoor attacks in DRL as it gives the adversary direct control over the agent – inducing predictable actions irrespective of the consequences or current state. Thus the adversary’s goal is to maximize the probability with which the agent chooses the target action $a^{+}$ given some poisoned state $\delta(s)$:
[…]
@@
The attack must also be stealthy, however, requiring the attacker to minimize the likelihood that the attack is detected while maximizing the chances that the agent is deployed in the real world. The most relevant definition of stealth can vary largely depending on the application domain, so in this work we will be using the most established and well defined notion of attack stealth in the literature (Rathbun et al., 2024; Kiourti et al., 2019) defined below:
[…]
@@
Here the adversary’s objective is to minimize the difference in value between a policy $\pi\sim\mathcal{L}(M)$ trained on the benign MDP, and a policy $\pi^{+}\sim\mathcal{L}(M^{\prime})$ trained on the adversarial MDP. In other words, the agent should still solve the benign MDP $M$ – making the victim less likely to detect any adversarial or unexpected behavior and more likely to deploy the agent in the real world.
3.2 Extended Problem Formulation via Training Time Detection
One key observation of this work is on the training-time detectability of prior attacks utilizing strategies focused on reward poisoning. These attacks, including but not limited to SleeperNets and TrojDRL, rely on arbitrarily large perturbations applied to the agent’s reward signal in order to solve both attack success and attack stealth. Specifically, SleeperNets and TrojDRL utilize unbounded and static reward poisoning strategies, respectively, as defined below given target action $a^{+}$:
Unbounded Reward $\begin{array}[]{cc}\textbf{Unbounded Reward Poisoning}&R^{u}(s_{p},a,s^{\prime})=\mathbbm{1}[a=a^{+}]-\gamma V_{\pi}^{M^{\prime}}(s^{\prime})\\ \textbf{Static Reward Poisoning}&R^{s}(s_{p},a,s^{\prime})=c\cdot(\mathbbm{1}[a=a^{+}]-\mathbbm{1}[a\neq a^{+}])\end{array}$ (3)
for some poisoned state $s_{p}\in S_{p}$. The reward poisoning induced by SleeperNets modifies the agent’s reward by $\gamma V_{\pi}^{M^{\prime}}(s^{\prime})$, effectively reducing their expected return to $\mathbbm{1}[a=a^{+}]$ in poisoned states. This term can grow arbitrarily large however, deviating significantly from the range of the benign reward function $R$. Similarly, the reward signal induced by static reward poisoning strategies scales linearly with the hyperparameter $c$, which often needs to be very large for attack success.
Figure 2: (Left) Rewards induced by SleeperNets and TrojDRL compared to benign behavior on the Highway Merge Environment. Perturbations detected by Equation 4 are marked with an alarm symbol. (Right) Performance of TrojDRL and SleeperNets before and after clipping is applied to their reward poisoning. After clipping we see a significant drop in Attack Success Rate.
In Figure 2 we give an example in the Highway Merge environment (Leurent, 2018). On the left we see that rewards obtained under TrojDRL and SleeperNets poisoning extend far beyond the range of the benign rewards, making both attacks detectable. This raises the question, how could a defender detect these perturbations in a principled manner? One property of reward functions the defender can leverage is that $R$ must be bounded by some finite $L,U\in\mathbb{R}$ such that $L=\inf[R]$ and $U=\sup[R]$. We assume the defender has sufficient knowledge of the task to compute $L$ and $U$ prior to training. From here they can define a strong yet simple set of detection rules. Let $\{r_{t}\}_{t=0}^{\infty}$ be a stream of rewards observed by the defender after reward perturbation. We then define their detection rule as:
$D(r_{t})\doteq\left\{\begin{array}[]{cc}\text{adversarial}&\text{if }r_{t}<L\vee r_{t}>U\\ \text{benign}&\text{otherwise}\end{array}\right.$ (4)
If the detection function $D$ ever returns the adversarial label, the victim will cease training and perform an investigation to remove the adversary. Under the scrutiny of this detector, the only remaining option for the attacker is to artificially clip their adversarial reward function to stay within $[L,U]$. The adversary won’t know these bounds a-priori, but can learn and update them as they observe the agent’s benign rewards. This greatly restricts their capabilities when only utilizing reward poisoning strategies, as seen in the right plot of Figure 2.
@@ 3.2 Extended Problem Formulation via Training Time Detection
To further demonstrate this lets consider a simple MDP as defined in Figure 3 with discount factor $\gamma$. Here the agent has two possible actions in the “Start” state, $a^{+}$ and $a$. When the agent takes action $a$ they prosper, receiving a reward of $+1$ on every time step for a discounted return of $\frac{\gamma}{1-\gamma}$ overall. When they take action $a^{+}$ they receive no reward and terminate immediately, receiving a discounted return of 00. In the “Prosper” state lets assume, for simplicity, that all agents always take action $a^{+}$. Now let’s consider the value of an arbitrary policy $\pi$ in the poisoned state $\delta(\text{Start})$ under unbounded, static, or bounded reward perturbations. Here bounded poisoning results in a reward of $1=\sup[R]$ if the target action $a^{+}$ is taken in a poisoned state or $0=\inf[R]$ otherwise.
Q-Values Unbounded Static Bounded
$Q_{\pi}^{M^{\prime}}(\delta(\text{Start}),a^{+})$ $1$ $c$ $1$
$Q_{\pi}^{M^{\prime}}(\delta(\text{Start}),a)$ 00 $\frac{\gamma}{1-\gamma}-c$ $\frac{\gamma}{1-\gamma}$
Table 2: Q values for an arbitrary policy $\pi$ in our example MDP under backdoor attacks utilizing unbounded, static, and bounded reward poisoning.
In Table 2 we can see the outcomes computed using Equation 3. If the agent chooses action $a$ under unbounded reward poisoning the term $-\gamma V_{\pi}^{M^{\prime}}(s^{\prime})$ will cancel out the value of the “Prosper” state, making $a$’s effective value 0, and making $a^{+}$ optimal. With static reward poisoning, and given sufficiently large $c$, the bonus reward of taking action $a^{+}$ can overwhelm the discounted return of the “Prosper” state, making $a^{+}$ optimal. However, when either of these attacks are subject to the detection function $D$ they can no longer induce such large reward perturbations. Thus, the largest reward for action $a^{+}$ they can give is $1$, and the lowest reward they can give for action $a$ is 0. Thus, for any $\gamma>\frac{1}{2}$ the value of $a$ will become larger than the value of $a^{+}$, making $a^{+}$ sub-optimal and resulting in the attack’s failure. Motivated by this investigation we add an additional constraint to our adversarial reward function $R^{\prime}$ with respect to the benign reward function $R$:
Reward Constraintssup[$\begin{array}[]{cc}\textbf{Reward Constraints}&\sup[R^{\prime}]\leq\sup[R]\text{ and }\inf[R^{\prime}]\geq\inf[R]\end{array}$ (5)

### v3

# Adversarial Inception Backdoor Attacks against Reinforcement Learning

@@ Abstract
Recent works have demonstrated the vulnerability of Deep Reinforcement Learning (DRL) algorithms against training-time, backdoor poisoning attacks. The objectives of these attacks are twofold: induce pre-determined, adversarial behavior in the agent upon observing a fixed trigger during deployment while allowing the agent to solve its intended task during training. Prior attacks assume arbitrary control over the agent’s rewards, inducing values far outside the environment’s natural constraints. This results in brittle attacks that fail once the proper reward constraints are enforced. Thus, in this work we propose a new class of backdoor attacks against DRL which are the first to achieve state of the art performance under strict reward constraints. These “inception” attacks manipulate the agent’s training data – inserting the trigger into prior observations and replacing high return actions with those of the targeted adversarial behavior. We formally define these attacks and prove they achieve both adversarial objectives against arbitrary Markov Decision Processes (MDP). Using this framework we devise an online inception attack which achieves an 100% attack success rate on multiple environments under constrained rewards while minimally impacting the agent’s task performance.
Machine Learning, ICML
[…]
@@ 1 Introduction
The wide-spread applicability of DRL (Mnih et al., 2013; Schulman et al., 2017) in security critical domains such as automated cyber defenses (Vyas et al., 2023), self-driving vehicles (Kiran et al., 2021), robotic warehouse management (Krnjaic et al., 2023), and space traffic coordination (Dolan et al., 2023) makes it a target for external adversaries wishing to influence the trained agent’s behavior. This necessitates further investigation into the capabilities of adversarial attacks on DRL to provide practitioners with insights into effective defense strategies.
In this work, we focus on gaining a deeper understanding of backdoor poisoning attacks, which manipulate an agent’s training to enable direct control over its behavior during deployment upon encountering a predefined trigger. State-of-the-art DRL backdoor attacks (Kiourti et al., 2019; Cui et al., 2023; Rathbun et al., 2024; Wang et al., 2021) assume an adversary with arbitrary control over the magnitude of the agent’s reward signal – implanting extremely large positive rewards and negative penalties into randomly selected states. This strongly biases the agent’s projected returns, causing the adversarial behavior to appear optimal whenever the trigger is observed.
However, RL environments have natural upper and lower bounds to their reward functions which these attacks ignore, exposing them to detection as outliers. Additionally, many practical RL implementations enforce reward post-processing such as clipping or normalization (Mnih et al., 2013; Huang et al., 2022) which can effectively erase perturbed rewards. These restrictions on the adversary prevents them from sufficiently biasing the agent’s projected returns, resulting in attack failure. Developing a principled poisoning strategy under these constraints poses a significant challenge as manipulating randomly selected states and rewards alone during training is no longer effective. Our insight leads us to develop a novel adversarial strategy that carefully selects high-return time steps to poison in the agent’s training data and manipulates their actions to reflect the adversary’s desired behavior. Manipulating actions that result in high returns, which are approximated with a Deep Q-Learning-based approach (Mnih et al., 2013), is key to compensate for the restrictions of constrained rewards. Through these strategies we provide multiple contributions towards our understanding of more realistic, reward constrained backdoor attacks against reinforcement learning. Specifically we:
@@
    Prove the limitations of prior poisoning attacks in achieving attack success against general DRL environments while respecting reward constraints.
@@
    Formulate the adversarial inception attack framework (Figure 1) which uses novel action manipulation techniques to guarantee backdoor attack success under constrained rewards while maintaining the victim agent’s performance in their intended task.
@@
    Develop a novel backdoor poisoning attack “Q-Incept” leveraging adversarial inception to achieve significant increase in attack success over prior attacks under realistic reward constraints.
@@
    Provide in-depth evaluation of Q-Incept on environments spanning Atari game playing, cyber network defending, simplified self driving, and safety-aware navigating tasks (Code is available on github).
Figure 1: Visualization of inception attacks. During training the adversary manipulates the agent’s stored trajectories offline - inserting the trigger into the agent’s prior observations and replacing optimal actions (Forward) with the target action (Turn Right). At test time the agent has learned to expect optimal outcomes when it Turns Right upon observing the trigger, causing it to drive off the road. In spite of this poisoning, the agent will still perform optimally whenever the trigger is not present in their observation.
@@ 2 Related Work and Background
When executing a backdoor attack, the adversary manipulates the Markov Decision Process (MDP) which an RL agent is being trained to optimize. MDPs are often defined as $M=(S,A,R,T,\gamma)$ where $S$ is the set of states in the environment, $A$ is the set of possible actions for the agent to take, $R:S\times A\times S\rightarrow\mathbb{R}$ is the reward function, $T:S\times A\times S\rightarrow[0,1]$ represents the transition probabilities between states given actions, and $\gamma\in[0,1]$ is the discount factor.
Backdoor attacks against DRL were first explored by Kiourti et al. (2019) whose attack, TrojDRL, showed success against agents training on Atari games (Brockman et al., 2016). Multiple other works (Wang et al., 2021; Yang et al., 2019; Yu et al., 2022; Cui et al., 2023) have used similar approaches in different domains – all statically altering the agent’s reward to a fixed ±$\pm c$, and many forcing the agent to explore the targeted action $a^{+}$ more often during training. Rathbun et al. (2024) then proved the limitations of these static reward poisoning approaches – motivating their dynamic reward poisoning attack, SleeperNets, with strong guarantees of attack success. Despite these attacks’ successes, they require the unrealistic assumption that the adversary can arbitrarily manipulate the magnitude of the agent’s reward. This assumption violates intrinsic constraints on rewards given in tasks such as standard Atari Benchmarks where rewards are constrained into a $[0,1]$ range (Mnih et al., 2013). Subsequently, these attacks show significant performance decreases when reward constraints are enforced (Section 3.2). Other works have studied different forms of training time attacks such as adversarial cheap talk (Lu et al., 2023) and policy replacement attacks (Rangi et al., 2022).
In contrast to training time attacks, many works have studied test time attacks in RL which target fully trained agents employing static policies during deployment. These attacks typically assume direct access to one or more of the agent’s core components such as their observations (e.g., sensors) (Lin et al., 2020), environment (Gleave et al., 2019), or actions (e.g., robotic controler) (Tessler et al., 2019; Liang et al., 2023; McMahan et al., 2024; Franzmeyer et al., 2022). Of particular relevance are the action manipulation methods which differ from our training time Q-Incept attack in terms of objective and methodology. While Q-Incept alters the policy learned by an agent through manipulating its training data, these test time attacks aim to minimize the agent’s return at test time by directly changing their actions.
@@ 3 Problem Formulation
In backdoor attacks against DRL there are two primary parties – the victim and the adversary. The victim attempts to train an agent to solve some benign MDP $M=(S,A,R,T,\gamma)$ where the state space $S\subset\mathbb{S}$ is a subset of some larger space (e.g., set of all possible 32x32 images) and the action space $A$ is discrete. The agent is trained with learning algorithm $\mathcal{L}(M)$ which returns a policy $\pi:\mathbb{S}\times A\rightarrow[0,1]$. The adversary induces the agent to instead train with respect to an adversarially modified MDP $M^{\prime}=(S\cup S_{p},A,R^{\prime},T^{\prime},\gamma,\delta,\beta)$ where $\delta:S\rightarrow\mathbb{S}$ applies the trigger to a given state, $S_{p}\doteq\{\delta(s)\;\forall s\in S\}$ is the set of poisoned states, $R^{\prime}$ is the adversarial reward function, $T^{\prime}$ is the adversarial transition function, and $\beta\in[0,1]$ bounds the frequency with which the agent can transition to poisoned states. Formally, $T^{\prime}$ and $R^{\prime}$ must respect the following constraints with respect to all $s,s^{\prime}\in S$ and $a\in A$:
$(1-\beta)T^{\prime}(s^{\prime}|a,s)+\beta T^{\prime}(\delta(s^{\prime})|a,s)=T(s^{\prime}|a,s)$ (1)
$R^{\prime}(s,a,s^{\prime})=R(s,a,s^{\prime})$ (2)
These constraints reflect the practical limitations of the attacker as they can not directly alter the source code of the environment simulator, nor can they modify the dynamics of real world environments. Instead, the attacker modifies the data which the agent is training on, inserting the trigger pattern $\delta$ into their state with probability $\beta$ after natural environment transitions have occurred. Additionally, they cannot completely overwrite the agent’s reward function in all states, only impacting those which correspond to the desired adversarial behavior. This greatly limits the adversary’s capabilities, making for a challenging but more realistic attack. Note that these constraints do not prohibit the adversary from altering transitions out of poisoned states $T(s^{\prime}|a,\delta(s))$ in theory, although in practice the adversary merely alters the agent’s training data. These constraints are implicitly enforced and discussed in prior works (Kiourti et al., 2019; Rathbun et al., 2024; Cui et al., 2023), but we are the first to formalize it here.
@@ 3.1 Adversarial Objectives
We consider “targeted attacks” (Kiourti et al., 2019) in which the desired adversarial behavior is a fixed target action $a^{+}\in A$. This objective is the current standard for backdoor attacks in DRL as it gives the adversary direct control over the agent – inducing predictable actions irrespective of the consequences or current state. Thus the adversary’s objective is to induce the agent to learn a poisoned policy $\pi^{+}\sim\mathcal{L}(M^{\prime})$ which takes action $a^{+}$ with high probability when observing the trigger:
Attack Success: $\begin{array}[]{cc}\textbf{Attack Success: }\max_{\pi^{+}}\frac{1}{|S|}[\sum_{s\in S}[\pi^{+}(\delta(s),a^{+})]]\end{array}$ (3)
@@
The attack must be stealthy, however, requiring the attacker to minimize the likelihood of detection by maintaining the agent’s return in benign states. The most relevant definition of stealth can vary depending on the application domain, so in this work we use the most established and well defined notion of attack stealth in the literature (Rathbun et al., 2024; Kiourti et al., 2019) defined below:
Stealth: $\begin{array}[]{cc}\textbf{Stealth: }\min_{\pi^{+}}[\frac{1}{|S|}\sum_{s\in S}[|V_{\pi^{+}}^{M}(s)-V_{\pi}^{M}(s)|]]\end{array}$ (4)
@@
where $\pi\sim\mathcal{L}(M)$, and $V_{\pi}^{M}(s)$, $V_{\pi^{+}}^{M}(s)$ are the expected values of policies $\pi$ and $\pi^{+}$ in MDP $M$ respectively given state $s$ (Sutton & Barto, 2018). Thus the adversary’s objective is to minimize the difference in value between an unpoisoned policy $\pi\sim\mathcal{L}(M)$, and a poisoned policy $\pi^{+}\sim\mathcal{L}(M^{\prime})$. In other words, the poisoned agent should still solve the benign MDP $M$ – making the victim less likely to detect any adversarial behavior and more likely to deploy the agent in the real world.
3.2 Why Reward Constrained Attacks?
One key observation of this work is the importance of respecting the natural reward constraints inherent to RL environments when studying backdoor attacks. Specifically, all RL environments have natural upper and lower bounds to their reward function in order to facilitate the convergence of cumulative returns. Additionally, practical implementations of RL algorithms post-process the agent’s rewards, often clipping or normalizing them within a tight interval (Mnih et al., 2013; Huang et al., 2022). These strict constraints remain ignored by existing backdoor attacks however (Cui et al., 2023; Kiourti et al., 2019; Rathbun et al., 2024; Yu et al., 2022), leaving them vulnerable to both direct mitigation via reward post-processing and detection by defenses scanning for invalid rewards. Specifically, SleeperNets and TrojDRL utilize dynamic ($R^{d}$) and static ($R^{s}$) reward poisoning strategies, respectively, as defined below given target action $a^{+}$:
$\begin{array}[]{c}R^{d}(\delta(s),a,s^{\prime})=\mathbbm{1}[a=a^{+}]-\gamma V_{\pi}^{M^{\prime}}(s^{\prime})\\ R^{s}(\delta(s),a,s^{\prime})=c\cdot(\mathbbm{1}[a=a^{+}]-\mathbbm{1}[a\neq a^{+}])\end{array}$ (5)
for some states $s,s^{\prime}\in S$. Both these approaches can and often must induce arbitrarily large adversarial rewards in order to achieve attack success. For instance, consider the example MDP defined in Figure 2 with discount factor $\gamma$. Here, the agent has two actions in the “Start” state, $a^{+}$ and $a$. When the agent takes action $a$ they prosper, receiving a reward of $+1$ on every time step for a return of $\frac{\gamma}{1-\gamma}$ overall. When they take action $a^{+}$ they receive no reward and terminate immediately, receiving a return of 00.
@@ 3.2 Why Reward Constrained Attacks?
Now let’s assume the MDP is impacted by a backdoor attack with target action $a^{+}$ using static or dynamic reward poisoning, as defined in Equation 5. Note that here the value of the “Prosper” state approaches infinity as $\gamma$ approaches $1$. Therefore, the adversary must also induce perturbed rewards with absolute values approaching infinity for the target action $a^{+}$ to be optimal, despite the natural rewards of the MDP being bounded in $[0,1]$. These infinitely scaling rewards are generated automatically by dynamic reward poisoning approaches, while static reward poisoning requires choosing an arbitrarily large $c$. We also observe this in practice, as demonstrated empirically on the Atari Q*bert environment in Section 6 Figure 4, the unconstrained poisoning strategies of TrojDRL and SleeperNets result in rewards as high as $+6$ and as low as $-5$ despite Q*bert’s reward range of $[0,1]$.
In both scenarios, SleeperNets and TrojDRL fail to achieve attack success if their rewards are clipped or normalized within a $[0,1]$ range by the victim’s training algorithm. Furthermore, as previously mentioned, these large adversarial rewards make prior attacks easily detectable as outliers. For instance, the simple, rule-based detector $D$ – defined below given input reward $r_{t}$ and benign reward function $R$ – detects both TrojDRL and SleeperNets in the prior examples with no false positives and without impacting agent performance.
$D(r_{t})\doteq\left\{\begin{array}[]{cc}\text{adversarial}&\text{if }r_{t}<\inf[R]\vee r_{t}>\sup[R]\\ \text{benign}&\text{otherwise}\end{array}\right.$ (6)
Therefore it is in the adversary’s best interest to ensure their rewards stay within the upper and lower bounds of $R$ to avoid detection. Motivated by the above exploration, we add an additional constraint to the adversary’s reward function $R^{\prime}$ with respect to the benign reward function $R$:
$\begin{array}[]{c}\sup[R^{\prime}]\leq\sup[R]\text{ and }\inf[R^{\prime}]\geq\inf[R]\end{array}$ (7)
Thus, the adversary must induce rewards no larger or smaller than those given in the benign MDP $M$. Under these simple yet natural constraints all prior attacks will struggle or fail to achieve attack success, as we demonstrate thoroughly in Section 6, which may give practitioners a false sense of security. Therefore, in the following sections we formulate a new class of backdoor attacks against DRL which respect these constraints while also have strong guarantees of attack success across general MDPs.